OpenStudy (anonymous):

If *line* RO=5x and RQ=12x-20, fins the value of x.

OpenStudy (anonymous):

5x= 12x-20 20=7x 20/7=7x/7 x=20/7

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

so you end up with 3 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

mmmk thank you

OpenStudy (anonymous):

2.85 to be exact

OpenStudy (anonymous):

but can i round it if i need to find what RO is? or do i just plugin 20/7

OpenStudy (anonymous):

you can round, no prob

OpenStudy (anonymous):

Thank you for your help. Me much happy now :P

OpenStudy (anonymous):

youre welcome